How many grams of Chlorine required to completely react with 0.650 g of Hydrogen to yield Hydrogen chloride? Also calculate the amount of HCl formed

Grade:11

1 Answers

Vanshika Govindani
2 years ago
H_{2} + Cl_{2}\rightarrow 2HCl
 
No. of moles required of H2 and Cl2 are equal to completely react to form HCl (2 moles).
So, moles of H2 = moles of Cl2
 
\frac{0.650}{2}= \frac{x}{70} 
where x is mass of required Cl2. x=22.75gm
 
now, moles of HCl formed= 2(moles of H2 or Cl2)
moles of HCl = 2(0.650/2)
\frac{y}{36.5}=0.650
 
y= amount of HCl formed =23.725gm